Direct Materials Variances

Silicone Engine Inc. produces wrist-worn tablet computers. The company uses Thin Film Crystal (TFC) LCD displays for its products. Each tablet uses one display. The company produced 490 tablets during December. However, due to LCD defects, the company actually used 510 LCD displays during December. Each display has a standard cost of $6.80. 510 LCD displays were purchased for December production at a cost of $2,885.

Determine the price variance, quantity variance, and total direct materials cost variance for December. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number. And, enter your final values to the nearest whole dollar.

Price variance $ Favorable
Quantity variance $ Unfavorable
Total direct materials cost variance $ Favorable
Answer #1

Price variance = (Standard price-actual price)actual quantity = (6.8*510-2885) = -583 Favorable

Quantity variance = (Standard qty-actual qty)Standard price = (490-510)*6.8 = 136 Unfavorable

Total direct material cost variance = Standard cost-actual cost = (490*6.8-2885) = -447 favorable